OpenStudy (en):

prove the formula: arccos x=1/2pi- arcsin x please i still dont get it

OpenStudy (emmatassone):

do you know a bit of calculus?

OpenStudy (en):

yep :) i'm solving bunch of problems .. i just dont get this one.. :/

OpenStudy (en):

sorry for the trouble :)

OpenStudy (emmatassone):

\[-\frac{ 1 }{ \sqrt{1-x²} }+\frac{ 1 }{ \sqrt {1-x²} }=0\] \[\int\limits_{}^{}-\frac{ 1 }{ \sqrt{1-x²} } dx+\int\limits_{}^{}\frac{ 1 }{ \sqrt {1-x²} }dx=0\] \[\arccos(x)+\arcsin(x)+\delta =0\] Where delta is a constant, evaluating in zero: \[\arccos(0)+\arcsin(0)+\delta=0\] \[\frac{ \pi }{ 2 } + \delta =0\] \[\delta = -\frac{ \pi }{ 2 }\] \[\arccos(x) + \arcsin(x) -\frac{ \pi }{ 2 }= 0\] Finally; \[\arccos(x)= \frac{ \pi }{ 2 }- \arcsin(x)\]

OpenStudy (emmatassone):

no problem ts not trouble xD

OpenStudy (emmatassone):

althought this demonstration is not general at all, if you look I choose arccos(0)=pi/2 but i could had chosen arccos(0)=3pi/2 ,5pi/2 , etc..

OpenStudy (en):

thanks :)))